OpenStudy (anonymous):

Differentiate f(x)=secx/ 1+secx I used quotient rule and so far I have F'(x)= secxtanx(1+secx) - (secx)(secxtanx) Step by step after this would be appreciated ! I thought I would multiply the secx and get Sec^2xtanx - sec^2xtanx Someone help please:(

OpenStudy (john_es):

\[f(x)=\frac{\sec x}{1+\sec x}=\frac{1/\cos x}{1+1/\cos x}=\frac{1}{1+\cos x}\\ f'(x)=\frac{1+\cos x+\sin x}{(1+\cos x)^2}\]

OpenStudy (anonymous):

The answer in the back of the book says secxtanx / (1+secx)^2 though

OpenStudy (anonymous):

I figured it out ! Wooo!

OpenStudy (anonymous):

Thank you for responding !!

OpenStudy (john_es):

Ok ;)